1) A magnifying glass has a convex lens with a focal length of 12 cm. A...

1)

A magnifying glass has a convex lens with a focal length of 12 cm. A rock is placed 30 cm away from the lens. What is the magnification factor of the rock?

2)

A laser has a wavelength of 1064 nm in air. It enters some diamond with an index of refraction of 2.42. What is the wavelength of the light in the diamond?

a. 530 nm

b. 2600 nm

c. 440 nm

d. 100 nm

e. 1060 nm
Answer #1

1.

focal length f = 12 cm

object distance u = 30cm

By lens formula

1/v - 1/u = 1/f

1/v = 1/f + 1/u

1/v = 1/12 +(1/-30)

v = 20cm

Magnification of lens is given by

M = v/u

M = 20cm / - 30cm

M = - 0.666

Forms a real image of the rock and is diminished.

magnification factor = - 0.66

2.

Wave length of laser in air \lambda air = 1064nm

Refractive index of diamond n = 2.42

Refractive index n = \lambda air / \lambda medium

n = 1064 nm / \lambda diamond

\lambdadiamond = 1064nm/ 2.42

= 439.67nm

= 440nm

The wavelength of light in diamond is 440nm